OpenStudy (anonymous):

ok, when adding numbers with like exponents, do you add or multiply the exponents? for example -2s^3 + 9s^3

OpenStudy (anonymous):

don't touch the exponents when adding or subtracting

OpenStudy (anonymous):

so that would be 7s^3?

OpenStudy (anonymous):

for example -2000 -9000 = 7000 by which i mean \[-2\times 10^3+9\times 10^3=7\times 10^3\]

OpenStudy (anonymous):

and so as you said \[-2s^3+9s^3=7x^3\]

OpenStudy (anonymous):

ok, great, thank you, will have another answer to verify here shortly lol

OpenStudy (anonymous):

not a problem of exonents . simply take it common